OpenStudy (anonymous):

plz i need help :( and you will get an medal Determine the width of the garden.Milo must put 50 feet of fencing around his rectangular garden. The length of the garden is one foot longer than twice the width.What is the solution set for 2w + 2(2w + 1) = 50 given the replacement set {6, 7, 8, 9}? A. 6 B. 7 C. 8 D. 9

OpenStudy (anonymous):

why don't they ask "solve for \(w\) in \( 2w + 2(2w + 1) = 50 \)"??

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

multiply out to get \[2w+4w+2=50\] then combine like terms on the left

OpenStudy (anonymous):

you should get \[6w+2=50\] solve by a) subtracting 2 then b) divide by 6